A bar magnet of moment 4Am is placed in a non uniform magnetic field. If the field strength at poles are 0.2T and 0.22T then the maximum couple acting on it is

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
As we know C = MB = 4* 0.22 = 0.88
 
Hope it helps
 
In case of any difficulty, please feel free to ask again.
 
Regards
Arun (askIITians forum expert)
Khimraj
4 years ago
C = M*B = 4* 0.22 = 0.88
..............................................................................